OpenStudy (anonymous):

Part A: Max rented a motorbike at $465 for 5 days. If he rents the same motorbike for a week, he has to pay a total rent of $625. Write an equation in the standard form to represent the total rent (y) that Max has to pay for renting the motorbike for x days. Part B: Write the equation obtained in Part A using function notation. Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals.

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

Do you have an idea on how to start?

OpenStudy (anonymous):

no idea. i only know how to do part c

OpenStudy (sleepyjess):

Ok, first we need to find the cost for 1 day

OpenStudy (anonymous):

which i cant do without part a and b

OpenStudy (sleepyjess):

wait, no

OpenStudy (anonymous):

one day is 93, right

OpenStudy (sleepyjess):

We need to find slope

OpenStudy (anonymous):

oh ok give me a sec

OpenStudy (sleepyjess):

but yes, 1 day is 93, wait, 1 day is the slope, duh

OpenStudy (anonymous):

so we have 5, 465 and 7, 625 625 - 465/7 - 5 160/2 80/1 ?????

OpenStudy (anonymous):

oooh

OpenStudy (sleepyjess):

now I'm confused...

OpenStudy (sleepyjess):

@ankit042

OpenStudy (anonymous):

it says write in standard for which is Ax + By = C, right?

OpenStudy (ankit042):

465=5m+c 625=7m+c now subtract two from one to get value of m

OpenStudy (anonymous):

i am so confused

OpenStudy (ankit042):

m will be the per day cost and c is upfront cost

OpenStudy (anonymous):

what is c? is it b in y = mx + b

OpenStudy (ankit042):

yeah I use y = mx+c notation

OpenStudy (anonymous):

oh

OpenStudy (ankit042):

To explain this let us assume we have this case where to rent a motor cycle you have to pay some upfront cost which is independent on number of day you hire the bike for (c) and m is the rent for each day so for x days one has to pay y dollars which can be represented as y=mx+c

OpenStudy (ankit042):

@sleepyjess still confused?

OpenStudy (sleepyjess):

uuuuummmm yeah

OpenStudy (anonymous):

i am supposed to write one equation to represent the total rent and x days do i do 465 = 5m + b or 625 = 7m + b

OpenStudy (ankit042):

using these equations you will find values of m and b then to represent it for x days you will say y=mx+b ( replacing values for m and b)

OpenStudy (ankit042):